# Progressive Quantum Algorithm for Maximum Independent Set with Quantum Alternating Operator Ansatz

### v2

@@ II.3 Review of QAOA+
III PROGRESSIVE QUANTUM ALGORITHM
Any arbitrary graph has at least one induced subgraph with the same MIS solution as it. Based on this property, PQA iteratively constructs a derived subgraph that includes the MIS solution of $G$ but is smaller in graph size. Naturally, solving the MIS problem on this subgraph requires fewer resources than directly solving it on the original graph. However, multiple derived subgraphs may exist, each with a different size but sharing the same MIS solution as $G$. As a result, various numbers of quantum gates and qubits are needed to implement the $p$-level QAOA+ ansatz for each subgraph. Fewer quantum gates and qubits are required in the realization of $e^{-i\beta_{i}H_{B}}$ and $e^{-i\gamma_{i}H_{C}}$ as the number of nodes and edges decreases. Especially, in the realization of $e^{-i\beta_{i}H_{B}}$, multiple $d(v)$-qubit controlled RX gates may be required. The overhead of implementing the $d(v)$-qubit controlled RX gate is costly but decreases when the node $v$ has fewer neighborhood nodes (i.e., a smaller $d(v)$) Vale et al. (2023); Tomesh et al. (2022b). To minimize the required qubits and reduce the overhead of implementing multi-qubit controlled RX gates, an ideal subgraph should not only have the same MIS solution as the target graph but also be small and sparse (i.e., with fewer edges and where the number of edges is less than the square of the number of nodes). Constructing an ideal subgraph is a key aspect of PQA. The following sections introduce the rules for constructing induced subgraphs and methods for estimating whether they share the same MIS solution as the target graph.
III.1 The construction rules of the subgraph
@@
To construct such an ideal subgraph, PQA starts from a sparse and small-sized initial subgraph $G_{0}$ and gradually expands it. During each expansion, the process favors adding the node with the lowest “closeness” to minimize the effect of the new node on the sparsity of the subgraph. The closeness is measured by a classical objective function
@@
where $V_{c}=V-V_{q}$ is the vertex candidate set, and $V_{q}$ is the vertex set of the current induced subgraph $G_{q}$. If the nodes $v_{c}$ and $v_{q}$ are adjacent (i.e., $(v_{c},v_{q})\in E$), $E(v_{c},v_{q})=1$. Otherwise, $E(v_{c},v_{q})=0$. The value of $C(v_{c})$ measures how the addition of vertex $v_{c}$ to the current subgraph $G_{q}$ affects its sparsity, and a small value of $C(v_{c})$ indicates a minimal change in sparsity. The choice of $G_{0}$ is not random, and it also involves the selection of the initial node in addition to graph expansion (i.e., the addition of new vertices and edges). The detailed selection rules are as follows.
@@
The choice of the first node. To construct a sparse initial subgraph $G_{0}$, the node with a minimum degree in the target graph $G$ is chosen as the first node of $G_{0}$. If there are $l(>1)$ nodes with a minimum degree, one is randomly chosen as the first node. The extension of subgraph. The addition of new nodes should have minimal impact on graph sparsity. That is, the node that has $\min C(v_{c})$ in the candidate set is the next vertex to join the current subgraph. If there are $l(>1)$ nodes with $\min C(v_{c})$, further calculations are performed. Specifically, for each node with $\min C(v_{c})$, we assume it is added to the current subgraph and calculate its corresponding $\min C(v_{c})$, resulting in $l$ calculation results. The node with the minimum value among these results is chosen to expand the current subgraph. If multiple nodes meet this criteria, one is randomly selected as the next node. During the extension, new edges may be introduced. Specifically, any edge $(v_{q},v_{new})$ should be added to the current subgraph if the newly added node $v_{new}$ is adjacent to existing nodes $v_{q}$, resulting in an expanded subgraph. Additionally, an example is provided in Appendix A to help readers understand the process of graph expansion.
@@ III.2 Progressive quantum optimization
After constructing an initial induced subgraph $G_{0}$, PQA builds a parameterized quantum circuit using the $p$-level QAOA+ ansatz to solve the MIS problem on $G_{0}$, with the parameters randomly initialized. After optimization, PQA continues to expand the graph size according to the given expansion rules and constructs $p$-level QAOA+ ansatz for the new subgraph to solve the MIS problem. At this stage, the PQC reuses the optimized QAOA parameters from the previous subgraph and re-optimizes these pre-trained parameters. This process is referred to as “parameter transfer” Galda et al. (2021); Shaydulin et al. (2023); Montanez-Barrera et al. (2024); Wurtz and Lykov (2021); Brandao et al. (2018); Li et al. (2023). After each optimization round, PQA gets an expectation function value $F_{G_{i},p}$ about the subgraph $G_{i}$, and PQA terminates if the values satisfy $\Delta_{1}\leq\xi$ and $\Delta_{2}\leq\xi$ or if the size matches that of the target graph in the worst case, where
@@
Here, $\xi$ is the tolerance for function values, and $F_{G_{i},p}$ ($F_{G_{i-1},p}$) represents the function value obtained using $p$-level QAOA+ ansatz to solve the MIS problem on the current (previous) subgraph $G_{i}$ ($G_{i-1}$). Here, we explain why $\Delta_{1}\leq\xi$ and $\Delta_{2}\leq\xi$ are used as the termination conditions. According to the properties of the MIS problem, adding a new vertex to the derived subgraph $G_{i}$ can cause the newly constructed subgraph $G_{i+1}$ to have an identical or larger independence number compared with $G_{i}$. In other words, $\beta(G_{i})\leq\beta(G_{i+1})$ (i.e., $F_{\text{max},G_{i}}\leq F_{\text{max},G_{i+1}}$). When PQA starts with a subgraph that has the same MIS solution as the target graph, the growth in the maximal expected function values $F_{\text{max},G_{i}}$ will cease as the subgraph grows (i.e., $F_{\text{max},G_{i}}$ reaches a steady value). Nevertheless, the value of $F_{\text{max},G_{i}}$ might also ascend with the increase of the graph size after a brief period of stagnation when PQA starts with a different initial subgraph. During the transient stationary phase, the addition of new nodes provides no benefit (i.e., without incrementing the maximal expectation function value). This occurs due to being trapped in a local subgraph, which is affected by sub-optimal node selection. For the latter case, the subsequent enlargement of the graph might enhance performance, but it concurrently requires more quantum resources to construct the corresponding quantum circuits. At this time, scaling the graph size without regard to resources is not advisable, and PQA should terminate the process in a timely manner. Given these considerations, we prioritize the stable phase of change in the expected function value during optimization, and PQA is halted if the expected function value remains stable over three consecutive optimization rounds. By this stage, we may have discovered a derived subgraph that includes the MIS solution of the target graph, obtaining an approximate or exact solution by solving the MIS problem on this subgraph. The quality of reused parameters can affect the optimization result for each subgraph. During actual optimization, the expectation function value may oscillate back and forth with increasing subgraph size when the parameter transfer performs poorly, delaying the fulfillment of the stop condition, such as shown in Appendix B. In such cases, PQA may not satisfy the termination condition and continue to execute optimization until reaching the same scale as the target graph, even though a subgraph with the same solution as the target graph has been constructed. This contradicts the original goal of PQA. To reduce running costs in PQA, we first aim to improve the quality of reused parameters by a greedy approach, as done in Ref. Shaydulin et al. (2019); Sack et al. (2023), when solving the MIS problem of $G_{0}$. Specifically, we can execute $c$ rounds of optimization when PQA solves the MIS problem on the initial induced subgraph, then select one set of the optimized parameters with the maximal function value in $c$ sets of results. The preserved parameters are reused to solve the MIS problem in the next subgraph. In addition, early exit mechanisms are required to end the low-cost optimization when appropriate. These mechanisms are based on evaluating the quality of the reused parameters and the optimization results for each subgraph. After each graph extension, we first decide whether to proceed with the optimization round by assessing the quality of the reused (initial) parameters. Generally, PQA can converge to quasi-optima after a few iterations (or even no optimization) when it starts with high-quality initial parameters. Otherwise, it may require abundant iterations to converge or get stuck in a low-quality local solution. Thus, we terminate PQA early and skip the current optimization round if the expected function value, based on the initial parameters, is significantly lower than that obtained from the previous subgraph’s optimization. If the quality of the initial parameters meets the rules we set, a new round of early exit judgments is carried out after the optimization. If the latest expected function value is significantly lower than the maximal value obtained from previous subgraphs, it indicates that the optimization has fallen into a poor local solution. At this time, PQA will halt further graph expansion and optimization, returning the maximal function value in this run and the corresponding output state. Otherwise, PQA continues to execute if it satisfies the above settings. In summary, to reduce the resources required during a PQA run, we introduced two early exit mechanisms that determine whether subsequent optimization rounds are necessary based on ‘flags’ (the quality of the initial parameters and the optimization results). Notably, PQA cannot get the optimal solution to the problem through an optimization run, and it simply aims to obtain a quasi-optimal solution with as few quantum resources as possible. More detailed procedures of PQA are shown in FIG. 4. PQA is a hybrid quantum-classical algorithm, where this concept encompasses two aspects. Firstly, when the QAOA+ algorithm is used to solve the MIS problem on a subgraph, the classical computer handles parameter optimization, while the quantum computer computes the expected function value. This constitutes an external parameter optimization loop, achieving synergy between classical and quantum computation. Secondly, in constructing the subgraph, the classical computer creates a new subgraph based on the structural information of the target graph. Meanwhile, the quantum computer solves the MIS problem on the subgraph and informs the classical computer whether to continue expanding the subgraph based on the optimization results. This forms another collaborative process. In summary, PQA leverages the parallel computing capabilities of the quantum computer to solve the MIS problem on each subgraph, while sharing part of the costs (such as obtaining target graph information and parameter adjustment) with the classical computer, rather than having either the quantum or classical computer handle all resource overheads alone.
IV RESULTS
In our work, the effects of graph construction rules and the initial graph size $n_{0}$ on the performance of PQA are first investigated in Section IV.1. Then, we compare the performance of different strategies on various graphs, with more details in Section IV.2.
IV.1 The effects of graph construction rules and $n_{0}$ on the performance of PQA
To investigate the impact of graph construction rules and the initial graph size $n_{0}$ on the performance of PQA, we randomly generate multiple ER graphs. For the given graph instances, we conduct $100\times p$ runs of PQA following various graph construction rules, where $n_{0}=1$, $\xi=0.1$, and the quantum state $|s\rangle$ is initialized as $|\mathbb{0}\rangle$ at each solving process. Subsequently, we calculate the achieved AR by PQA at level $p=1$ under different construction rules. Notably, the AR is the ratio between the expectation function value obtained by PQA on the final induced subgraph and $F_{\text{max},G}$. Here, $F_{\text{max},G}$ denotes the maximal expectation function value on the target graph, equivalent to the independence number of the target graph. The AR is less than or equal to one, and an AR of one signifies that PQA achieves the MIS solution of the target graph by solving the MIS problem on the current subgraph.
IV.1.1 The effects of construction rules on the performance of PQA.
For each graph instance, we first calculate the optimal and average approximation ratio (OAR, AAR) obtained by multiple optimization runs of PQA when it constructs subgraphs according to different rules. Then, we average the obtained OAR and AAR across all graph instances to get the overall average OAR and AAR for PQA. From these simulation results given in FIG. 5, we observe that the expansion rules play a more significant role in the performance of PQA compared with the choice of the first node. Specifically, PQA achieves a much higher AR when following our expansion rules instead of random expansion. Given these phenomena, we analyze the variation of the expected function value during optimization and find that the parameter transfer tends to perform poorly (i.e., getting stuck in a low-quality local solution) when random vertex additions cause significant changes in the graph structure, leading to early termination of the optimization run. This may occur because significant changes in the graph structure lead to drastic shifts in the function landscape, rendering the transferred parameters suboptimal for the new landscape. In addition, FIG. 6 provides the total number of iterations required by PQA under various construction rules while achieving their respective OAR at $p=1$. These findings suggest that PQA requires far fewer iterations to obtain a quasi-optimal solution when following our graph expansion rules, with a $95\%$ reduction in iterations compared with random first node and expansion (i.e., ’NN’). In summary, the results presented in FIG. 5 and FIG. 6 collectively demonstrate the efficiency of our proposed graph construction rules.
Figure 5: The effects of graph construction rules on the AR achieved by PQA, where ‘NN’ suggests PQA arbitrarily chooses the first node and subsequent nodes without any rules. ‘NY’ represents that the first node is randomly chosen, while other newly added nodes obey our predetermined expansion rules.
Figure 6: The effects of graph construction rules on the total consumption of iterations to get the quasi-optimal solution.
Figure 7: The OAR achieved by PQA at various level depths $p$ and $n_{0}$ on given ER graphs, where $n_{0}$ is the number of nodes in the initial induced subgraph $G_{0}$. Each solid dot in the curve represents the AR obtained by PQA at level depth $p$ when it starts from a subgraph with $n_{0}$. Specifically, the curve reflects the AR obtained by DS-QAOA+ under various level depths when $n_{0}=14$.
IV.1.2 The effects of $n_{0}$ on the performance of PQA.
Furthermore, we explore the impact of $n_{0}$ on the performance of PQA under various $p$. In our simulations, we vary the level depth from one to five and perform $100\times p$ runs of PQA at each level. The resulting AR versus the level depth and $n_{0}$ is depicted in FIG. 7. These simulation results suggest that PQA may require more level depths to get the exact solution when the initial problem size is large. Concretely, PQA can achieve the exact MIS solution (i.e., OAR = 1) at $p=1$ when starting with initial subgraphs of size $n_{0}=2,4,6$, while for PQA with $n_{0}=14$ (i.e., DS-QAOA+), its required level depth is at least five. In general, the algorithm often requires multiple optimization runs at specific level depths to obtain the exact solution or reach a given OAR threshold. During parameter optimization, each run consists of numerous iterations, with each iteration requiring multiple calls to the quantum processing units (QPUs) to compute the expectation function value. More iterations increase the demand on QPUs. To comprehensively examine the effect of $n_{0}$ on the performance of PQA, we analyze the total number of optimization runs (TRs) and iterations (TITRs) when PQA initializes from various $n_{0}$ values and achieves a given OAR at the required level depths (RLDs). Specifically, for each graph instance, we first calculate the OAR achieved by multiple runs at each level depth to determine the required level depths for each algorithm to reach the given OAR. Then, we calculate the number of runs $Q$ when the resulting AR exceeds the given OAR threshold over multiple optimization runs. We regard the ratio of $Q$ to the total number of runs $R$ at level depth $p$ as the probability of exceeding that threshold. The reciprocal of this probability gives the TRs required for the algorithm to reach the given OAR. In addition, we calculate the average iterations (AITRs) needed for an optimization run at the RLDs, and the TITRs required for the algorithm to achieve the given OAR at the RLDs equals the product of TRs and AITRs. For each graph instance, we conduct a similar analysis to obtain the RLDs, TRs, and TITRs. For each graph type, we compute the average of multiple RLDs, TRs, and TITRs to estimate the number of level depths, total optimization runs, and total iterations required for the algorithm to achieve a given OAR when solving the MIS problem on a given graph type.
Table 2: The required level depths (RLDs), total number of optimization runs (TRs), total iterations (TITRs), total qubits (T-qubits), total circuit depths (TCDs), total multi-qubit CR (T-MQCR) gates, and total quantum (T-Q) gates for PQA to achieve the given OAR under different values of $n_{0}$.
 OAR $n_{0}$ RLDs TRs TITRs T-qubits TCDs T-MQCR gates T-Q gates
0.9 2 1 1.5562 76.0375 8.7926 23.568 6.6096 17.5852
4 3.2 8.4233 704.2204 74.54 1079.1239 340.6935 716.1204
6 1.6 1.4397 95.0968 11.5176 66.9053 21.2897 43.2461
8 2.4 1.9594 125.2674 20.1942 152.0068 49.0538 98.1076
10 3.4 4.1255 294.5122 49.5827 489.149 158.6944 317.3888
12 4.2 7.1945 549.2766 100.3954 1507.5039 490.7771 981.5542
14 4.8 46.8667 2387.0363 656.13332 14169.9319 4613.4662 9226.9324
0.85 2 1 1.5562 76.0375 8.7926 23.568 6.6096 17.5852
4 3 9.7833 832.7083 94.7 1252.4839 397.3335 829.4004
6 1.4 1.5901 98.982 12.7205 67.6558 21.5299 43.7264
8 2.2 1.9921 125.2955 20.8618 150.6104 48.6473 97.2946
10 3.2 4.783 371.4308 60.6338 584.0127 189.7596 379.5192
12 4.0 5.5698 426.9543 77.7713 1125.6186 366.4565 732.913
14 4.6 33.0571 1521.2784 462.8 7244.8851 2358.7998 4717.5996
Moreover, for each graph instance, the total number of qubits (T-qubits), total circuit depths (TCDs), total multi-qubit controlled RX (T-MQCR) gates, and total quantum (T-Q) gates demanded by each algorithm are also calculated when they achieve the given OAR. Combining with these indicators, we aim to clearly reflect the resource expenses for PQA with different $n_{0}$. The circuit depth measures how many “layers” of quantum gates are executed in parallel, and all quantum gates that can be applied simultaneously in the circuit are regarded as one layer. The circuit depth reflects the execution time of the quantum calculation process Ni et al. (2024); Guerreschi and Matsuura (2019). Deeper circuits require more time for execution, increasing the risk of noise and errors Du et al. (2022). In this work, the circuit depth is obtained by directly using the DAGCircuit interface in MindSpore Quantum Developer (2021). The gate resource refers to the number and types of quantum gates required in a quantum circuit, such as single-qubit rotation gates and multi-qubit controlled rotation (CR) gates, with the latter being more complex to implement on quantum hardware. Gate resources reflect the operational complexity of quantum circuits, with more gates generally leading to greater complexity. In our work, we count the number of single-qubit RX gates, single-qubit RZ gates, and multi-qubit controlled RX gates. The sum of these gives the consumption of quantum gates required by the PQC with the $p$-level QAOA+ ansatz. In our calculations, the number of $d(v)$-qubit controlled RX gates in one level QAOA+ ansatz is incremented by one if $d(v)\geq 1$. The total quantum resources are the product of the average consumption per run and TRs at their respective RLDs. From these simulation results shown in Table 2, we observe that PQA’s resource demands increase as it starts with a larger initial induced subgraph. Specifically, PQA with $n_{0}=2$ achieves the given OAR while consuming only about $4.718\%$ of the quantum resources required by DS-QAOA+ (i.e., PQA with $n_{0}=14$). This represents a savings of approximately $95.9\%$ of iterations compared with DS-QAOA+, revealing substantial efficiency gains with PQA. In conclusion, these above investigations emphasize the sensitivity of PQA to the choice of $n_{0}$. Its performance may potentially decline when initialized with a large-scale initial subgraph. To maintain robust performance in subsequent simulations, we opt for $n_{0}=2$. This decision is informed by our observation that smaller values of $n_{0}$ tend to yield more reliable results, ensuring the stability and effectiveness of PQA across varied scenarios.

### v3

@@ II.2 Review of QAOA
II.3 Review of QAOA+
The original QAOA searches for the ground state of $H_{C}$ from a Hilbert space that includes both feasible and infeasible quantum states, the output states may not always satisfy the independence constraint and violate one or more constraints constraints_QAOA ; QIRO . To overcome this challenge, Hadfield et al. graph_coloring1 ; alternating_QAOA_ansatz introduce a novel quantum alternating operator ansatz algorithm denoted as QAOA+. Different from incorporating the independence constraint into the target Hamiltonian utilizing the Lagrange multiplier, QAOA+ encodes constraints into the mixer Hamiltonian $H_{B}$. This allows QAOA+ to search for the optimal solution within a subspace that only contains feasible quantum states. Consequently, the target Hamiltonian
$H_{C}=\sum_{u=0}^{n-1}\frac{\sigma_{u}^{z}-I}{2}$ (8)
is obtained by converting the $x_{u}$ in $-C_{1}(x)$ to a quantum spin $\frac{I-\sigma_{u}^{z}}{2}$, thereby encoding the MIS solution into the ground state of $H_{C}$. Similar to QAOA, QAOA+ approaches the ground state of $H_{C}$ by alternately applying target and mixer unitaries on an initial quantum state $|s\rangle$ that is required to be trivial to implement. However, in QAOA+, the initial quantum state must be the superposition of feasible states. This requirement ensures that the feasible subspace contains no states that violate the constraints. In addition, the initial quantum state $|s\rangle$ may not be the ground state of $H_{B}$ in QAOA+, unlike in QAOA.
Figure 3: The implementation of $e^{-i\beta_{i}B_{u}}$ when $d(u)\geq 1$. The multi-qubit controlled $R_{x}$ gate can be decomposed into two multi-qubit controlled gates and a single-qubit controlled $R_{x}$ gate, utilizing an auxiliary qubit gate_decomposition . The auxiliary qubit is initialized as $|0\rangle$ and is flipped to $|1\rangle$ only when the quantum states of all adjacent vertices are $|0\rangle$ (i.e., $f(u)=0$). Once this condition is met, an $R_{x}$ gate is applied on the $u$-th qubit. After the operation, the state of the auxiliary qubit can be reset from $|1\rangle$ to $|0\rangle$ using the right multi-qubit controlled gate operation. Otherwise, the auxiliary qubit is always $|0\rangle$ when $f(u)>0$, and there is an Identity operation applied to the $u$-th qubit.
The design of the mixing Hamiltonian $H_{B}$ must preserve the feasibility of the subspace (i.e., ensure the subspace only has quantum states obeying constraints) and explore a new feasible subspace (i.e., provide transitions between various feasible states) alternating_QAOA_ansatz . For an independent vertex subset $V_{s}$, deleting an arbitrary vertice $v_{a}$ (i.e., the state of $x_{v_{a}}$ is converted from 1 to 0) from $V_{s}$ does not change its independence. However, a new vertex $v_{a}$ can be added to $V_{s}$ (i.e., flipping $x_{v_{a}}$ from 0 to 1) without destroying its independence only when there are no adjacent vertices of $v_{a}$ in $V_{s}$. Thus, for any node, controlling its bit-flip operation based on the states of its neighbors is sufficient to both add and remove vertices while preserving the independence property alternating_QAOA_ansatz . The bit-flip of $v_{a}$ occurs only if
$f(v_{a})=\sum_{j=1}^{d(v_{a})}x_{v_{j,a}}=0,$ (9)
where $v_{j,a}$ is the $j$-th adjacent vertex of $v_{a}$, and $d(v_{a})$ is the number of adjacent vertices of $v_{a}$, with $j=1,\cdots,d(v_{a})$. Mapping to a quantum circuit, for any vertex $u$, a Pauli-X operation is applied on the $u$-th qubit (i.e., this qubit is flipped) if and only if all quantum states of adjacent vertices of $u$ are $|0\rangle$ (i.e., $f(u)=0$). Otherwise, a Pauli-I operation is applied on the $u$-th qubit. For the node $u$, its corresponding partial mixer Hamiltonian $B_{u}$ can be described as
The mixer Hamiltonian is given by $H_{B}=\sum_{u=0}^{n-1}B_{u}$ alternating_QAOA_ansatz . Specifically, the quantum state resulting from applying $e^{-i\beta_{i}B_{u}}$ to the single feasible quantum state $|x_{o}\rangle$ is
$e^{-i\beta_{i}B_{u}}|x_{o}\rangle=\begin{cases}&R_{x}(2\beta_{i})|x_{o}\rangle,f(u)=0\\ &|x_{o}\rangle,f(u)>0,\end{cases}$ (11)
where the detailed result of $R_{x}(2\beta_{i})|x_{o}\rangle$ is $\cos 2\beta_{i}|x_{o}\rangle-i\sin 2\beta_{i}|x_{0}\cdots\bar{x_{u}}\cdots x_{n-1}\rangle$, and $\bar{x_{u}}$ denotes the state of $u$-th qubit is converted from $|0\rangle$ to $|1\rangle$ (or ($|1\rangle$ to $|0\rangle$), implementing the addition or removal of node $u$ without generating infeasible quantum states. The circuit implementation of $e^{-i\beta_{i}B_{u}}$ is shown in FIG. 3 when the node has $d(u)\geq 1$ adjacent vertices. Additionally, Table 1 is provided to illustrate clearly the differences between using QAOA and QAOA+ to solve the MIS problem.
III Progressive quantum algotirhm
Any arbitrary graph has at least one induced subgraph that preserves the same MIS solution as the original graph. Inspired by this property, PQA aims to iteratively construct a derived subgraph that is likely to contain the MIS solution of $G$ while being smaller in size. Naturally, solving the MIS problem on this subgraph requires fewer qubits than directly solving it on the original graph. To construct such an induced subgraph, PQA iteratively expands the subgraph using heuristic expansion strategies designed for this purpose, starting from a sparse and small subgraph. Once the subgraph has been expanded, PQA proceeds to tackle the MIS problem within this subgraph by utilizing QAOA+. Subsequently, PQA iterates the processes of graph expansion and problem-solving until a previously specified termination condition is fulfilled. In PQA, the purpose of the graph expansion is to improve the probability that the constructed subgraph contains the MIS solution of the original target graph, while solving the MIS problem on each subgraph using QAOA+ serves to estimate whether the current subgraph contains the MIS solution of the original graph by obtaining the expected function values. The expansion strategies and the setting of the stopping condition will be detailed in the following.
III.1 The graph construction and expansion strategies
@@
The original graph may have multiple derived subgraphs, each differing in size but retaining the same MIS solution. The overhead of implementing the multi-qubit controlled $R_{x}$ gate is expensive and increases polynomially as the number of control qubits (that correspond to the neighborhood nodes) grows gate_decomposition ; gate_decomposition_MIS . To minimize the required qubits and reduce the overhead of implementing multi-qubit controlled $R_{x}$ gates, an ideal subgraph should not only have the same MIS solution as the target graph but also be small and sparse. Here, the small property helps reduce the number of qubits required, and the sparse property reduces the overheads associated with multi-qubit controlled gates. To construct such an ideal subgraph, PQA starts from a sparse and small-sized initial subgraph $G_{0}$ and gradually expands it. During each expansion, the process favors adding the node with the lowest “closeness” to minimize the effect of the new node on the sparsity of the subgraph. The closeness is measured by a classical objective function
@@
where $V_{c}=V-V_{q}$ is the vertex candidate set, and $V_{q}$ is the vertex set of the current induced subgraph $G_{q}$. If the nodes $v_{c}$ and $v_{q}$ are adjacent (i.e., $(v_{c},v_{q})\in E$), $E(v_{c},v_{q})=1$. Otherwise, $E(v_{c},v_{q})=0$. The value of $C(v_{c})$ measures how the addition of vertex $v_{c}$ to the current subgraph $G_{q}$ affects its sparsity, and a small value of $C(v_{c})$ indicates a minimal change in sparsity.
@@
By starting with a small subgraph and gradually expanding its size, PQA can effectively control the growth of qubits, enhance computational efficiency, and minimize unnecessary resource usage. The construction of $G_{0}$ is not random, and it also involves the selection of the initial node and graph expansion (i.e., the addition of new vertices and edges). The detailed strategies are as follows. Additionally, an example is provided in Appendix A to help readers understand the process of graph expansion. The choice of the first node. To construct a sparse initial subgraph $G_{0}$, the node with a minimum degree in the target graph $G$ is chosen as the first node of $G_{0}$. If there are $l(>1)$ nodes with a minimum degree, one is randomly chosen as the first node. The extension of subgraph. The addition of new nodes should have minimal impact on graph sparsity. That is, the node that has $\min C(v_{c})$ in the candidate set is the next vertex to join the current subgraph. If there are $l(>1)$ nodes with $\min C(v_{c})$, further calculations are performed. Specifically, for each node with $\min C(v_{c})$, we assume it is added to the current subgraph and calculate its corresponding $\min C(v_{c})$, resulting in $l$ calculation results. The node with the minimum value among these results is chosen to expand the current subgraph. If multiple nodes meet this criteria, one is randomly selected as the next node. During the extension, new edges may be introduced. Specifically, any edge $(v_{q},v_{new})$ should be added to the current subgraph if the newly added node $v_{new}$ is adjacent to existing nodes $v_{q}$, resulting in an expanded subgraph.
@@ III.2 Progressive quantum optimization
After constructing an initial induced subgraph $G_{0}$, PQA builds a PQC using the $p$-level QAOA+ ansatz to solve the MIS problem on $G_{0}$, with the parameters randomly initialized. After optimization, PQA continues to expand the graph size according to the given expansion strategies and constructs $p$-level QAOA+ ansatz for the new subgraph to solve the MIS problem. At this stage, the PQC reuses the optimized QAOA parameters from the previous subgraph and re-optimizes these pre-trained parameters. This process is referred to as “parameter transfer” PIL ; parameter_transferability ; parameter_transfer ; transfer_learning2024 ; angle_conjecture ; leapfrogging . After each optimization round, PQA gets an expectation function value $F_{G_{i},p}$ about the subgraph $G_{i}$, and PQA terminates if the values satisfy $\Delta_{1}\leq\xi$ and $\Delta_{2}\leq\xi$ or if the size matches that of the target graph in the worst case, where
@@
Here, $\xi$ is the tolerance for function values, and $F_{G_{i},p}$ ($F_{G_{i-1},p}$) represents the function value obtained using $p$-level QAOA+ ansatz to solve the MIS problem on the current (previous) subgraph $G_{i}$ ($G_{i-1}$). Here, we explain why $\Delta_{1}\leq\xi$ and $\Delta_{2}\leq\xi$ are used as the termination conditions. According to the properties of the MIS problem, adding a new vertex to the derived subgraph $G_{i}$ can cause the newly constructed subgraph $G_{i+1}$ to have an identical or larger independence number compared with $G_{i}$. In other words, $\beta(G_{i})\leq\beta(G_{i+1})$ (i.e., $F_{\text{max},G_{i}}\leq F_{\text{max},G_{i+1}}$). When PQA starts with a subgraph that has the same MIS solution as the target graph, the growth in the maximal expected function values $F_{\text{max},G_{i}}$ will cease as the subgraph grows (i.e., $F_{\text{max},G_{i}}$ reaches a steady value). Nevertheless, the value of $F_{\text{max},G_{i}}$ might also ascend with the increase of the graph size after a brief period of stagnation when PQA starts with a different initial subgraph. During the transient stationary phase, the addition of new nodes provides no benefit (i.e., without incrementing the maximal expectation function value). This occurs due to being trapped in a local subgraph, which is affected by sub-optimal node selection. For the latter case, the subsequent enlargement of the graph might enhance performance, but it concurrently requires more quantum resources to construct the corresponding quantum circuits. At this time, scaling the graph size without regard to resources is not advisable, and PQA should terminate the process in a timely manner. Given these considerations, we prioritize the stable phase of change in the expected function value during optimization, and PQA is halted if the expected function value remains stable over three consecutive optimization rounds. By this stage, we may have discovered a derived subgraph that includes the MIS solution of the target graph, obtaining an approximate or exact solution by solving the MIS problem on this subgraph. During actual optimization, we find the expectation function value may oscillate back and forth with increasing subgraph size. In such cases, PQA may not satisfy the termination condition and continue to execute optimization until reaching a similar scale to the target graph, even though a subgraph with the same solution as the target graph has been constructed. This contradicts the original goal of PQA. To avoid unnecessary waste of resources in the optimization run, we set up early exit mechanisms that are required to end the low-cost optimization when appropriate. These mechanisms are based on evaluating the quality of the reused parameters and the optimization results for each subgraph. After each graph extension, we first decide whether to proceed with the optimization round by assessing the quality of the reused (initial) parameters. Generally, PQA can converge to quasi-optima after a few iterations (or even no optimization) when it starts with high-quality initial parameters. Otherwise, it may require abundant iterations to converge or get stuck in a low-quality local solution. Thus, we terminate PQA early and skip the current optimization round if the expected function value, based on the initial parameters, is significantly lower than that obtained from the previous optimization. If the quality of the initial parameters meets the condition we set, a new round of early exit judgments is carried out after the optimization. If the latest expected function value is significantly lower than the maximal value obtained from previous subgraphs, it indicates that the optimization has fallen into a poor local solution. At this time, PQA will halt further graph expansion and optimization, returning the maximal function value in this run and the corresponding output state. Otherwise, PQA continues to execute if it satisfies the above settings. In summary, to reduce the optimization resources required during a PQA run, we introduced two early exit mechanisms that determine whether subsequent optimization rounds are necessary based on ‘flags’ (the quality of the initial parameters and the optimization results). More detailed procedures of PQA are shown in FIG. 4, which may help readers more intuitively understand the PQA process. PQA is a hybrid quantum-classical algorithm that integrates quantum and classical computing to solve optimization problems, combining the strengths of both computational paradigms. Firstly, when the QAOA+ algorithm is used to solve the MIS problem on a subgraph, the classical computer is responsible for optimizing the parameters, while the quantum computer calculates the expected function value. This creates a feedback loop where classical and quantum components collaborate, with the classical computer adjusting parameters based on the output of the quantum computation. Secondly, the classical computer generates a new subgraph based on the structural information of the target graph. The quantum computer then solves the MIS problem on this subgraph and communicates the optimization results to the classical computer, which decides whether to expand the subgraph further based on the feedback. This forms another layer of collaboration between the classical and quantum components. In summary, PQA harnesses the superposition and parallelism of the quantum computer to solve the MIS problem on each subgraph. It collaborates with the classical computer, which handles tasks such as obtaining target graph information and adjusting parameters. This approach ensures that both quantum and classical components share the resource overheads rather than having one handle the entire burden alone.
IV Numerical simulations
In our simulations, we compare the performance among various algorithms on three graph types, including ER graphs with edge probabilities of 0.4 and 0.5, as well as 3-regular graphs. There are 20 randomly generated graphs in each type, with 14 nodes (i.e., $n=14$). For PQA, the tolerance on function values is set to $\xi=0.1$, and the quantum state $|s\rangle$ is initialized to $|\mathbb{0}\rangle$ at the start of each optimization round. For both DS-QAOA+ and QLS, the initial quantum state is set to $|\mathbb{0}\rangle$ at the beginning of each optimization run. For QLS, the local subgraph is determined by a root node, $n_{\text{root}}$, and a free parameter, $N_{s}$. Specifically, the subgraph consists of the root node and all nodes whose distance from the root is $\leq N_{s}$, along with all relevant edges among these nodes. Here, the root node is randomly selected, and the free parameter $N_{s}$ is the minimum value in the list $l_{\text{max}}$. The list $l_{\text{max}}$ stores the maximum shortest distances from any node to all other nodes in the original graph. For the three algorithms, the classical outer parameter optimization loop ends when the convergence tolerance MLI satisfies $\varepsilon\leq 0.001$. Additional details on the comparisons are provided below.
IV.1 The AR obtained by various algorithms with the same number of optimization runs and level depths
To compare the performance of various quantum algorithms, we present the optimal AR (OAR) and average AR (AAR) achieved by each algorithm on different types of graphs, using the same number of runs and level depths. Concretely, we set the level depth from one to five, and each algorithm is executed with $100$ optimization runs at each level depth. For each graph type, this procedure yields 20 independent OAR (and AAR) values. The mean OAR (AAR) is then calculated as the average of these 20 values. Additionally, to establish a baseline for comparison, we include the AR obtained by the hill-climbing algorithm (i.e., a classical local search algorithm with wide applications) hill_climbing and a classical counterpart to PQA, which follows the same graph construction and expansion strategies described in Section III.1. However, instead of employing QAOA+ to approximate the MIS solution for each subgraph, we solve each subproblem using the hill-climbing algorithm. Notably, when solving each subproblem using the hill-climbing algorithm, the initial solution is inherited from the approximate solution obtained from the previous subgraph, except for the first subgraph, where an initial feasible solution is randomly generated.